2) You are given the following information about an insurers book of business: . Policyholders are divided in three groups: low-risk, medium-risk, and of policyholders are considered medium-risk and 50% are considered high-risk. low-risk. . The probability that a low-risk policyholder has filed a claim is 0.05 .The probability that a randomly selected policyholder is considered medium- risk and has filed a claim is 0.035. . Given a policyholder is considered high risk, the probability that the pol- icyholder has filed a claim is equal to 4 times the probability that a randomly selected policyholder is considered high-risk. Calculate the probability a polieyholder is considered low-trak, gven they have filed a claim. A)0.10 B)0.17 C)0.21 D) 0.26 E)0.34 The circled answer is wrong, please show clear steps to arrive to a correct answer I will rate more points.
Answer #1

P(filed a claim)=P(medium risk)*P(claim|medium risk)+P(low risk)*P(claim|low risk)+P(high risk)*P(claim|high risk)=0.035+0.5*0.05+0.15*4*0.15=0.15

therefore P(low risk|claim)=P(low risk)*P(claim|low risk)/P(claim)=0.5*0.05/0.15=0.17

option B